jack had $69 and wilson had $21 at first. they donated a sum of money to charity and paid the amount equally. after that, Jack had 4 times as much money left as wilson. how much was the sum of money donated?
Answer by striker1997(2)

take the sum both of them donated to be x, after both of them donated x amount, jack has 4 times the amount wilson has. So we can form this equation:4(21-x)=69-x
as jack has 4 times the amount wilson has, wilson has to multiply his amount left by 4 so that they will have the same amount.
4(21-x) is wilson's amount left.
(69-x) is jack's amount left.
so now we multiply 4 into (21-x) and that gives u (84-4x)
now put the x on the left and the amount on the right, so it gives u -4x+x=69-84.
after u calculate, it becomes -3x=-15.
so x=5.